PLEASE HELP WITH THIS WORD PROBLEM.....THE SECOND ANGLE OF A TRIANLGE IS 30 DEGREES LESS THAN THE FIRST ANGLR. THE THIRD ANGLE IS TWICE THE SECOND ANGLE. HOW LARGE ARE THE ANGLES?
Answer by stanbon(57213)

THE SECOND ANGLE OF A TRIANLGE IS 30 DEGREES LESS THAN THE FIRST ANGLR. THE THIRD ANGLE IS TWICE THE SECOND ANGLE. HOW LARGE ARE THE ANGLES?
-------------------
Let 1st angle be "x":
Then 2nd angle be "x-30"
And 3rd angle is 2(x-30) = 2x-60
------------------------------------
Equation:
x + x-30 + 2x-60 = 180
4x -90 = 180
4x = 270
x = 67.5 degrees
------------------------
x-30 = 37.5 degrees
------------------------
2x-60 = 75 degrees
==============================
Cheers,
Stan H.
==============================